OpenStudy (anonymous):

ind the perimeter of an object which has a length of √180 feet and a width of √125 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. What type of object might this be?

OpenStudy (anonymous):

Find*

OpenStudy (anonymous):

length =sqrt(5*5*5)= 5sqrt(5) breadth=sqrt(6*6*5)= 6sqrt(5) thus perimeter =2(length+breadth)=2(5sqrt(5)+6sqrt(5))=22sqrt(5)